[figure: 9.7 cm]

QUANTUM WELL EFFECTS IN MAGNETORESISTANCE OF

ULTRA-THIN LAYERD STRUCTURES OF

FERROMAGNETIC/NONMAGNETIC SEMICONDUCTORS

J. Barnaś^{a,b}, S. Krompiewski^{b}, H.W. Kunert^{c}

aDepartment of Physics, A.M. University, Umultowska 85, 61-614 Poznań, Poland bInstitute of Molecular Physics PAS, Smoluchowskiego 17, 60-179 Poznań, Poland

c Department of Physics, University of Pretoria, 001 Pretoria, South Africa

In this paper we analyse theoretically the spin-valve magnetoresistance in a sandwich structure consisting of two thin films of a ferromagnetic semiconductor, which are separated by an ultra-thin layer of a nonmagnetic semiconductor. A simple parabolic band is assumed to describe electronic properties of the semiconductors (spin-split band in the ferromagnetic case). Electronic conductivity is calculated within the Boltzmann kinetic approach. Quasi-two dimensional electronic subband structure and quantum well states due to size quantization are taken into account. A quantity of particular interest is the relative change of the resistance when the magnetic configuration changes from antiparallel to parallel one, and the dependence of this ratio on the thickness of the nonmagnetic and ferromagnetic sublayers. Considerations are limited to current flowing in the plane of the structure. We found strong exponential decrease of the magnetoresistance with increasing thickness of the nonmagnetic spacer, and a more complex dependence on the thickness of the magnetic films. Variation of the magnetoresistance with the angle between magnetic moments of the ferromagnetic films is also studied.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : J. Barnaś

Address for correspondence : J. Barnaś Department of Physics ul. Umultowska 85 61-614 Poznań, Poland

Email address : barnas@main.amu.edu.pl